HUBBLE SPACE TELESCOPE DAILY REPORT #3046 PERIOD COVERED: 0000Z (UTC) 01/30/02 - 0000Z (UTC) 01/31/02 Daily Status Report as of 031/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD/MA1 9137 (Quasar Absorbers and Large Scale Structure) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to perform spectroscopy of 15 bright quasars in a 22 square degree region that has well-sampled galaxy redshifts. No problems were encountered. 1.2 Completed Four Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.3 Completed Seven Sets of WF/PC-2 9318 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform the generic target version of the WFPC2 Archival Pure Parallel program. The program was used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The re-acquisition failure mentioned in 2.1 and HSTAR 8487 occurred for the fifth iteration of this proposal and affected two observations when the take data flag remained down. Otherwise, there were no further reported problems. 1.4 Completed STIS/CCD 8901 (Dark Monitor-Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. There was no anomalous activity. 1.5 Completed Two Sets of STIS/CCD 9164 (Helical Flows and Rotation in Protostellar Jets) The Space Telescope Imaging Spectrograph (CCD) was used to study the velocity structure of protostellar jets transverse to the jet flow axis. HST resolution would allow the search for rotation, e.g. a helical flow pattern, around the axis of an optical jet. Such a detection would provide strong evidence that jets act as outflowing channels for angular momentum, thus allowing infalling material to fully accrete onto the protostar. There were no problems. 1.6 Completed WF/PC-2 9319 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to execute a POMS test proposal, designed to simulate future scientific plans. The proposal completed without incident. 1.7 Completed Two Sets of WF/PC-2 8941 (Cycle 10 UV Earthflats) The WF/PC-2 was used to monitor flat field stability by obtaining sequences of earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set. The proposal had no problems. 1.8 Completed STIS/CCD 8903 (Bias Monitor - Part 1) The Space Telescope Imaging Spectrograph (CCD) was to be used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 to build up high-S/N superbiases and track the evolution of hot columns. The re-acquisition failure mentioned in 2.1 and HSTAR 8487 occurred for this proposal and affected the observation when the take data flag remained down. 1.9 Completed STIS/CCD 9088 (Next Generation Spectral Library of Stars) The Space Telescope Imaging Spectrograph (CCD) was used to produce a "Next Generation'' Spectral Library of 600 stars for use in modeling the integrated light of galaxies and clusters by using the low dispersion UV and optical gratings of STIS. The library will be roughly equally divided among four metallicities, very low {Fe/H < -1.5}, low {-1.5 < Fe/H < -0.5}, near-solar {-0.5 < Fe/H < 0.1}, and super-solar {Fe/H > 0.1}, well-sampling the entire HR-diagram in each bin. Such a library will surpass all extant compilations and have lasting archival value, well into the Next Generation Space Telescope era. No problems occurred. 1.10 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) The Space Telescope Imaging Spectrograph (MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise in order to monitor the effects of thermal changes on the NUV dark rate. No problems occurred. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 4 Successful: 4 Scheduled Re-acquisitions: 11 Successful: 10 Per HSTAR 8485, a re-acquisition at 030/123153Z succeeded on the second attempt following a scan step limit being exceeded on FGS-2 in the first attempt. No observations were affected. HSTAR 8486 was written when another re-acquisition at 030/172417Z succeeded on the second attempt following a scan step limit being exceeded on FGS-2 in the first attempt. Again, no observations were affected. HSTAR 8487 describes an re-acquisition failure at 030/203701Z due to the scan step limit being exceeded on FGS-2. The following acquisition was successful. Proposals described in 1.3 and 1.8 were affected. 2.2 FHST Updates: Scheduled: 10 Successful: 10 2.3 Operations Notes: Using ROP SR-1A, the SSR-1 EDAC error counter was cleared three times. A TTR was written when a TDRS-W software update problem at 030/1759Z caused events and data to be lost. Accordingly, at 030/1734Z and at 030/2000Z, the engineering recorder was commanded in order to recover information. This exercise utilized ROP SR-8A. Another TTR was generated when there was a required re-transmit at 031/0829Z during a 486 load. The engineering status buffer limit for SESBSLD was updated at 031/0829Z as directed by ROP DF-18A. A target-of-opportunity is being inserted into the current SMS. The observations are scheduled to begin tomorrow morning. Processing of a new SMS is now underway. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW